OpenStudy (anonymous):

How many times as intense as a standard earthquake is a earthquake measuring 3.1 on the richter scale?

OpenStudy (anonymous):

what is the intensidy of a standard earthquake?

hero (hero):

I'm pretty sure you have the formula for this somewhere.

OpenStudy (anonymous):

i havent been given any formuals other than basic ones

OpenStudy (anonymous):

R = log[ I ÷ I0 ]

OpenStudy (anonymous):

M=log I/S

jimthompson5910 (jim_thompson5910):

Hint: Solve for I (capital i)

jimthompson5910 (jim_thompson5910):

After that, plug in M = 3.1

jimthompson5910 (jim_thompson5910):

Keep in mind that \[\Large I_{0}\] is the "standard earthquake"

OpenStudy (anonymous):

so if i solve for I what do i start with?

jimthompson5910 (jim_thompson5910):

well the intensity of it anyway

jimthompson5910 (jim_thompson5910):

I would convert the given equation to exponential form

jimthompson5910 (jim_thompson5910):

to start things off

OpenStudy (anonymous):

i was gunna start with m=log 3.1/s

jimthompson5910 (jim_thompson5910):

no, m = 3.1 since m is the magnitude

OpenStudy (anonymous):

3.1=log I/S

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now convert to exponential form

OpenStudy (anonymous):

i dont know how to do that

jimthompson5910 (jim_thompson5910):

Use the idea that \[\Large \log_{b}(x) = y\] is the same as \[\Large b^y = x\]

OpenStudy (anonymous):

i still dont understand im left with log I/S correct

OpenStudy (anonymous):

and log is to the base 10

jimthompson5910 (jim_thompson5910):

When you convert \[\Large 3.1 = \log\left(\frac{I}{S}\right)\] to logarithmic form, you get \[\Large 10^{3.1} = \frac{I}{S}\]

jimthompson5910 (jim_thompson5910):

Now solve for i

OpenStudy (anonymous):

10^3.1 becomes 1258

jimthompson5910 (jim_thompson5910):

1258.92541179417 but you pretty much have it

OpenStudy (anonymous):

ok so do i bring over the I or what?

jimthompson5910 (jim_thompson5910):

\[\Large 3.1 = \log\left(\frac{I}{S}\right)\] \[\Large 10^{3.1} = \frac{I}{S}\] \[\Large 1258.92541 = \frac{I}{S}\]

jimthompson5910 (jim_thompson5910):

now isolate 'i'

jimthompson5910 (jim_thompson5910):

by moving that 's' over

OpenStudy (anonymous):

so it becomes 1258/S

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

Multiply both sides by S.

OpenStudy (anonymous):

1258s ?= I

jimthompson5910 (jim_thompson5910):

yes \[\Large 3.1 = \log\left(\frac{I}{S}\right)\] \[\Large 10^{3.1} = \frac{I}{S}\] \[\Large 1258.92541 = \frac{I}{S}\] \[\Large 1258.92541S = I\] \[\Large I = 1258.92541S\]

jimthompson5910 (jim_thompson5910):

The right side says "1258.92541 times S"

OpenStudy (anonymous):

ok so final answer is the earthquake is 1258 time S for intense than the standard earthquake?

jimthompson5910 (jim_thompson5910):

no, just roughly 1258.92541

jimthompson5910 (jim_thompson5910):

a magnitude 3.1 is roughly 1258.92541 times more intense than a standard earthquake

jimthompson5910 (jim_thompson5910):

S is the intensity of the standard earthquake

OpenStudy (anonymous):

ok thank you a bunch

jimthompson5910 (jim_thompson5910):

np